6x-4+3/2x=5x+11
I have put x all on one side 6x+3/2x-5x=11+4
I made everything with x an fraction 12/2x+3/2x-10/2x=15
12+3-5=10
10/2=5
5x=15
x=3

Answer by jim_thompson5910(35256)

You made a mistake in red below

12+3-5=10

when it *should* be

12+3-10 = 5

which becomes 5/2

So

12/2x+3/2x-10/2x=15

5/2x=15

5x = 15*2

5x = 30

x = 30/5

x = 6

So the solution is x = 6